# A BENCHMARK AND EVALUATION FOR REAL-WORLD OUT-OF-DISTRIBUTION DETECTION USING VISION-LANGUAGE MODELS

### v2

@@ 2.2 ImageNet-FS-X
Background. An existing problem setting for OOD detection with covariate shifts is full-spectrum OOD (FS-OOD) [16, 22]. In FS-OOD, the ID data is constructed by adding data with the same semantics but different covariate distributions (covariate-shifted ID) to the training data (training ID). When using ImageNet-1k as the training ID, variant datasets (e.g., ImageNet-V2 [26], ImageNet-R [27], and ImageNet-C [28]) are used as covariate-shifted IDs (ImageNet-FS). Covariate-shifted ID is added only during evaluation to assess the performance change of ID data due to covariate shift. However, differences in covariate distributions between ID and OOD data are not considered, making it difficult to systematically evaluate semantic and covariate shifts separately.
@@ 2.3 Wilds-FS-X
Wilds consists of datasets spanning 10 different domains. For this benchmark, we utilized: iWildCam and FMoW. iWildCam is a dataset of wildlife photos taken by camera traps. Semantic shift corresponds to different animal species, and covariate shift corresponds to changes in camera locations of camera photos (Fig. 2). The labels comprise 182 classes, including 181 animal species, and an additional class labeled “empty” for images containing no animals. Among these, 91 animal labels were randomly selected as ID data, while the remaining 90 animal labels, along with the “empty” class, were used as OOD data. FMoW is a RGB satellite image dataset. Semantic shift corresponds to the building or land use categories, and covariate shift corresponds to changes in the year when the images were captured (Fig. 2). The dataset includes 62 facility types. From these, 31 classes were randomly chosen as ID data, while the remaining 31 classes were designated as OOD data. The specific sample sizes for the splits in this experiment are shown in Table 1. Using these datasets, we can evaluate the detection of fine-grained semantic shifts and robustness to covariate shifts required in real-world scenarios.
@@ 4 Conclusion
Appendix A OOD Score examples
@@ A.2 ImageNet-FS-X
Appendix B Covariate shifts analysis

### v3

@@ 2.2 ImageNet-FS-X
Background. An existing problem setting for OOD detection with covariate shifts is full-spectrum OOD (FS-OOD) [16, 22]. In FS-OOD, the ID data is constructed by adding data with the same semantics but different covariate distributions (covariate-shifted ID) to the training data (training ID). When using ImageNet-1k as the training ID, variant datasets (e.g., ImageNet-V2 [26], ImageNet-R [27], and ImageNet-C [28]) are used as covariate-shifted IDs (ImageNet-FS). Covariate-shifted ID is added only during evaluation to assess the performance change of ID data due to covariate shift. However, differences in covariate distributions between ID and OOD data are not considered, making it difficult to systematically evaluate semantic and covariate shifts separately. Moreover,  [29] proposed a benchmark that allows flexible evaluation of semantic shift and covariate shift in OOD data and analyzed their respective impacts. However, this does not account for covariate shifts in ID data and does not cover the FS-OOD problem setting.
@@ 2.3 Wilds-FS-X
Wilds consists of datasets spanning 10 different domains. For this benchmark, we utilized: iWildCam and FMoW.
iWildCam is a dataset of wildlife photos taken by camera traps. Semantic shift corresponds to different animal species, and covariate shift corresponds to changes in camera locations of camera photos (Fig. 2). The labels comprise 182 classes, including 181 animal species, and an additional class labeled “empty” for images containing no animals. Among these, 91 animal labels were randomly selected as ID data, while the remaining 90 animal labels, along with the “empty” class, were used as OOD data. FMoW is a RGB satellite image dataset. Semantic shift corresponds to the building or land use categories, and covariate shift corresponds to changes in the year when the images were captured (Fig. 2). The dataset includes 62 facility types. From these, 31 classes were randomly chosen as ID data, while the remaining 31 classes were designated as OOD data. The specific sample sizes for the splits in this experiment are shown in Table 1. Using these datasets, we can evaluate the detection of fine-grained semantic shifts and robustness to covariate shifts required in real-world scenarios.
@@ 4 Conclusion
ACKNOWLEDGEMENT
This work was partially supported by JST JPMJCR22U4 and JSPS 25H01164.
@@ ACKNOWLEDGEMENT
Appendix A Experimental Detail
In this section, we describe the details of each CLIP-based out-of-distribution (OOD) detection method employed in our experiments.
A.1 Zero-shot OOD detection
MCM and GL-MCM. We utilized the pre-trained CLIP-ViT-B/16 model (https://github.com/openai/CLIP). The same pre-trained CLIP model was consistently used across the other methods as well.
NegLabel. For the list of additional labels, we adopted the label set provided by previous studies.
EOE. This method uses a large language model (LLM) to generate additional labels. We employed gpt-3.5-turbo-16k for this purpose. Following prior work, which suggests designing LLM prompts based on visual similarity to produce suitable outlier labels, we used two types of prompts: the Far OOD prompt, which generates domain-level labels for Common-OOD evaluation; and the Near OOD prompt, which generates fine-grained labels for Hard-OOD and our proposed benchmark.
CLIPN. We employed the “no” Text Encoder provided by previous work as a pre-trained model. Among the two proposed OOD scoring algorithms—the Competing-to-win Algorithm and the Agreeing-to-differ Algorithm—we chose the latter (CLIPN-A), as it has been reported to deliver superior detection performance.
A.2 Few-shot OOD detection
The results for each few-shot method were averaged over 3 random seeds.
CoOp and LoCoOp. We used the pre-trained CLIP-ViT-B/16 model and trained with 16 shots, following the same hyperparameters as in prior work (e.g.training epochs = 50, learning rate = 0.002, batch size = 32, token length = 16).
NegPrompt. Prompt learning was conducted in two stages: first, positive prompts were trained for 100 epochs; then, with the positive prompts frozen, negative prompts were trained for 10 epochs. The loss weights ($\beta$ = 0.1, $\gamma$ = 0.05) were taken from previous work. Training was performed with 16 shots. Since the learning rate and batch size were not explicitly stated, we set them to 0.002 and 32, respectively.
IDPrompt. Due to its higher training cost, training was conducted with only 1 shot, as in prior studies. Each image was randomly cropped into 256 patterns, and the top 32 were used for training. Other hyperparameters (e.g.training epochs = 3, learning rate = 0.005, batch size = 1) also followed previous implementations.
Appendix B OOD Score examples
@@ B.2 ImageNet-FS-X
Appendix C Covariate shifts analysis